# Schrödinger Bridge Matching for Tree-Structured Costs and Entropic Wasserstein Barycentres

### v1

@@ B.2 Characterisation of TreeSB Solution
[…]
@@
We now show that for a Markov reference process ${\mathbb{Q}}$, the ($\mathrm{TreeSB}_{\mathrm{dyn}}$) solution is also Markov. Following Léonard [2013], to present the results we will use the following characterisation of Markovianity for a path measure $\mathbb{P}$ on the tree structure. Consider an edge $e\in\mathcal{E}$ and a time $t_{e}\in[0,T^{e}]$. Consider the time $t_{e}$ as splitting the continuous tree into two distinct sections (note that such a splitting may not be unique, as the chosen point may correspond to a vertex; in such cases consider any such split into two distinct parts). Denote the restrictions of a process $X$ to these two parts as $X_{\leq}$ and $X_{\geq}$. Then we say $\mathbb{P}$ is Markov to mean that $\mathbb{P}(X_{\leq}\in\cdot,X_{\geq}\in\cdot\cdot|X_{t}^{e})=\mathbb{P}(X_{\leq}\in\cdot|X_{t}^{e})\mathbb{P}(X_{\geq}\in\cdot\cdot|X_{t}^{e})$ for any such split (together with the technical assumption that some time-marginal of $\mathbb{P}$ is $\sigma$-finite; see discussion in Léonard [2013]). Note that the Brownian reference process considered in the main paper is Markov.
@@
Proof.
The proof follows that of Proposition 2.10 in Léonard [2013]. We outline the following changes to the notation for our setting, then the argument follows the same way.
Similarly to Léonard [2013] we now make the following claim, from which the result follows.
Claim B.6.
Fix a time on the tree $t_{e}$ as described above. Fix a $z\in{\mathbb{R}}^{d}$, a measure $\mu\in\mathcal{P}({\mathbb{R}}^{d})$, and path measures on the ‘before’ and ‘after’ sections ~$\tilde{\mathbb{P}}^{t_{e},z}_{\leq}\in\mathcal{P}(\mathcal{C}_{\mathcal{T}}^{\leq}\cap\{X_{t_{e}}=z\})$ and ~$\tilde{\mathbb{P}}^{t_{e},z}_{\geq}\in\mathcal{P}(\mathcal{C}_{\mathcal{T}}^{\geq}\cap\{X_{t_{e}}=z\})$. Consider minimising ${\mathrm{KL}}(\cdot\|{\mathbb{Q}})$ over path measures $\mathbb{P}\in\mathcal{P}(\mathcal{C}_{\mathcal{T}})$ constrained to satisfy $\mathbb{P}_{t_{e}}=\mu$, $\mathbb{P}^{t_{e},z}_{\leq}=\tilde{\mathbb{P}}^{t_{e},z}_{\leq}$, and $\mathbb{P}^{t_{e},z}_{\geq}=\tilde{\mathbb{P}}^{t_{e},z}_{\geq}$. Then the objective ${\mathrm{KL}}(\cdot\|{\mathbb{Q}})$ attains its unique minimum at $\mathbb{P}^{*}(\cdot)=\int_{{\mathbb{R}}^{d}}\tilde{\mathbb{P}}^{t_{e},z}_{\leq}\otimes\tilde{\mathbb{P}}^{t_{e},z}_{\geq}\mu({\mathrm{d}}z)$.
Given the claim, the result follows according to the following argument: Suppose for a contradiction that the SB solution, here denoted $\tilde{\mathbb{P}}$, was not Markov. Then, there exists some time $t_{e}$ and a correspond split of the tree such that ~${\tilde{\mathbb{P}}(\cdot|X_{t_{e}}^{e})\neq\tilde{\mathbb{P}}_{\leq}(\cdot|X_{t_{e}}^{e})\otimes\tilde{\mathbb{P}}_{\geq}(\cdot|X_{t_{e}}^{e})}$. Applying the above claim with $\mu=\tilde{\mathbb{P}}_{t_{e}}$, we see that $\mathbb{P}^{*}$ and $\tilde{\mathbb{P}}$ have the same marginals at all time-points on the tree, but $\mathbb{P}^{*}$ attains a strictly lower KL divergence ${\mathrm{KL}}(\mathbb{P}^{*}\|{\mathbb{Q}})<{\mathrm{KL}}(\tilde{\mathbb{P}}\|{\mathbb{Q}})$, contradicting the optimality of $\tilde{\mathbb{P}}$.
The proof of the claim uses Jensen’s inequality and proceeds exactly as the proof of Claim 2.11 in Léonard [2013], with the appropriate notation changes. ∎
@@
As we are considering a Brownian reference process, the assumptions in question are that $\int\lVert x\rVert^{2}{\mathrm{d}}\mu_{i}(x)<\infty$, and ${\mathrm{H}}(\mu_{i})<\infty$ for each $i\in\mathcal{S}$. One can then verify (as in De Bortoli et al. [2024], Lemma D.2) that the criteria in Proposition B.3 hold by taking functions $A$ and $B$ to be quadratic, from which uniqueness and existence of the solution follow. From Proposition B.4 we have that the solution is in $\mathcal{R}_{\mathcal{S}}({\mathbb{Q}})$, and from Proposition B.5 we have that it is Markov.
[…]
The remainder follows the standard argument characterising the SB solution using the decomposition according to potentials (see e.g. Nutz [2021]). Consider static couplings in the constraint set $\Pi\in\Gamma(\{\mu_{i}\}_{i\in\mathcal{S}}):=\{\Pi\in\mathcal{P}(({\mathbb{R}}^{d})^{|\mathcal{V}|}:\Pi_{i}=\mu_{i}~{}~{}\forall i\in\mathcal{S}\}$ such that ${\mathrm{KL}}(\Pi|{\mathbb{Q}}_{\mathcal{V}})<\infty$. By the above, we have that ${\mathbb{E}}_{\Pi}[\log(\tfrac{{\mathrm{d}}\Pi^{0}}{{\mathrm{d}}{\mathbb{Q}}_{\mathcal{V}}})]=\sum_{i\in\mathcal{S}}\int\log f_{i}{\mathrm{d}}\mu_{i}$, which is in particular independent of the choice of $\Pi$ (for a precise statement taking care regarding the integrability of the potentials, follow the argument of Proposition 2.17 in Nutz [2021]). Therefore, for any such $\Pi$ we have
@@ B.3 Properties of the Tree-Structured Projections
We now move on to proving properties of the TreeIMF procedure. We follow the presentation of Shi et al. [2023]. We begin by proving the properties of the reciprocal and Markovian projections defined in Definitions 3.3 and 3.4. For a full set of required assumptions, see the assumptions A.1, A.2 and A.3 in Shi et al. [2023] Appendix C.2, which are standard in the literature and we assume to hold along each edge.
@@
${\mathrm{KL}}(\mathbb{P}\|\tilde{\mathbb{P}})$ $={\mathrm{KL}}(\mathbb{P}_{r}\|\tilde{\mathbb{P}}_{r})+{\mathbb{E}}_{X_{r}\sim\mathbb{P}_{r}}\big{[}{\mathrm{KL}}\big{(}\mathbb{P}(\cdot|X_{r})\|\tilde{\mathbb{P}}(\cdot|X_{r})\big{)}\big{]}$ (31)
@@
$(\displaystyle{\mathrm{KL}}\big{(}$$\mathbb{P}(\cdot|X^{e_{k}},...,X^{e_{1}},X_{r})\|\tilde{\mathbb{P}}(\cdot|X^{e_{k}},...,X^{e_{1}},X_{r})\big{)}$
@@
${\mathrm{KL}}(\mathbb{P}\|\tilde{\mathbb{P}})$ $=\sum_{k=1}^{|\mathcal{E}|}{\mathbb{E}}_{\begin{subarray}{c}X_{e_{k-1}}\sim\mathbb{P}^{e_{k-1}}(\cdot|X^{e_{k-2}},...,X^{e_{1}},X_{r})\\ \vdots\\ X_{e_{1}}\sim\mathbb{P}^{e_{1}}(\cdot|X_{r})\\ X_{r}\sim\mathbb{P}_{r}\end{subarray}}\Big{[}{\mathrm{KL}}\big{(}\mathbb{P}^{e_{k}}(\cdot|X_{s(e_{k})})\|\tilde{\mathbb{P}}^{e_{k}}(\cdot|X_{s(e_{k})})\big{)}\Big{]}$ (36)
@@
We now consider the Markov projection defined in Definition 3.4. Note that the restriction of the TreeSB solution to each edge is itself an SB (because it is Markov and reciprocal), and thus can be associated with an SDE [Dai Pra, 1991]. In the definition of the Markov class in Definition 3.1 we define the Markov class $\mathcal{M}_{\mathcal{T}}$ via considering SDEs with locally Lipschitz drifts. The restriction to locally Lipschitz drifts is a technical requirement for applying the entropic version of Girsanov’s theorem; this is standard in the literature and does not affect our methodology. Following Shi et al. [2023], we now provide a result showing that the Markov projection also solves a minimisation problem.
@@
We now analyse the individual KL expressions ${\mathrm{KL}}(\Pi^{e}(\cdot|X_{u})\|{\mathbb{M}}^{e}(\cdot|X_{u}))$ along each edge $e=(u,v)$, using the proof techniques of Proposition 2 in Shi et al. [2023].
In particular, applying the argument in the proof of Proposition 2 in Shi et al. [2023], one sees that each conditional process $\Pi_{\cdot|0}^{e}$ is Markov and can be associated with $(X_{t}^{e})_{t\in[0,T^{e}]}$ given by
@@
Therefore, letting the restriction of the Markov process to edge $e$ (denoted above as ${\mathbb{M}}^{e}$) be associated with a process ${\mathrm{d}}Y_{t}^{e}=v_{e}(t,Y_{t}^{e}){\mathrm{d}}t+\sigma{\mathrm{d}}B_{t}^{e}$ such that ${\mathrm{KL}}(\Pi^{e}(\cdot|X_{u})\|{\mathbb{M}}^{e}(\cdot|X_{u}))<\infty$, with $v_{e}$ locally Lipschitz, then (using e.g. Léonard [2012], Theorem 2.3) one obtains
@@
We also note that, as an instance of bridge matching, along each edge the process $\Pi_{t}^{e}$ and its corresponding Markovian projection ${\mathbb{M}}_{t}^{e,*}$ satisfy the same Fokker-Planck equation (Peluchetti [2022], Theorem 2). Thus by the uniqueness of the solutions of the Fokker-Planck equations under A.1 and A.3 in Shi et al. [2023] (see e.g. Bogachev et al. [2021]), they share the same marginals ${\mathbb{M}}_{t}^{e,*}=\Pi_{t}^{e}$. ∎
@@ B.4 TreeIMF Convergence
We follow the presentation of convergence in Shi et al. [2023], but with the appropriate modifications to the proofs for the tree-based setting.
Lemma B.7 (Pythagorean property, compare to Shi et al. [2023], Lemma 6).
@@
Using these expressions, by applying the same algebraic manipulations as the proof of Lemma 6 in Shi et al. [2023] to each term in the summations, one obtains
@@
Proof of (45): The second part also follows similarly to Shi et al. [2023], but instead conditioning on the values at $\mathcal{S}$. Let $\Pi^{*}=\operatorname{proj}_{\mathcal{R}_{\mathcal{S}}({\mathbb{Q}})}(\mathbb{P})=\mathbb{P}_{\mathcal{S}}{\mathbb{Q}}_{\cdot|\mathcal{S}}$. Using the change of measure formula for KL divergence and the fact that $\Pi$ and $\Pi^{*}$ have the same bridges conditional on $\mathcal{S}$,
@@
In light of the Pythagorean property in Lemma B.7, this follows from the same compactness argument as in Proposition 7 and Theorem 8 in Shi et al. [2023]. ∎

### v2

@@ B.2 Characterisation of TreeSB solution
The solution $\mathbb{P}^{SB}$ to ($\mathrm{TreeSB}_{\mathrm{dyn}}$) (if it exists) is in the reciprocal class $\mathcal{R}_{\mathcal{S}}({\mathbb{Q}})$, i.e. $\mathbb{P}^{SB}=\mathbb{P}^{SB}_{\mathcal{S}}{\mathbb{Q}}_{\cdot|\mathcal{S}}$.
@@
We now show that for a Markov reference process $\mathbb{Q}$}, the ($\mathrm{TreeSB}_{\mathrm{dyn}}$) solution is also Markov. Following [60], to present the results we will use the following characterisation of Markovianity for a path measure $\mathbb{P}$ on the tree structure. Consider an edge $e\in\mathcal{E}$ and a time $t_{e}\in[0,T^{e}]$. Consider the time $t_{e}$ as splitting the continuous tree into two distinct sections (note that such a splitting may not be unique, as the chosen point may correspond to a vertex; in such cases consider any such split into two distinct parts). Denote the restrictions of a process $X$ to these two parts as $X_{\leq}$ and $X_{\geq}$. Then we say $\mathbb{P}$ is Markov to mean that $\mathbb{P}(X_{\leq}\in\cdot,X_{\geq}\in\cdot\cdot|X_{t}^{e})=\mathbb{P}(X_{\leq}\in\cdot|X_{t}^{e})\mathbb{P}(X_{\geq}\in\cdot\cdot|X_{t}^{e})$ for any such split (together with the technical assumption that some time-marginal of $\mathbb{P}$ is $\sigma$-finite; see discussion in [60]). Note that the Brownian reference process considered in the main paper is Markov.
@@
For Markov reference measure $\mathbb{Q}$}, the solution $\mathbb{P}^{SB}$ to ($\mathrm{TreeSB}_{\mathrm{dyn}}$) (if it exists) is Markov.
Proof.
The proof follows that of Proposition 2.10 in [60]. We outline the following changes to the notation for our setting, then the argument follows the same way.
Similarly to [60] we now make the following claim, from which the result follows.
Claim B.6.
Fix a time on the tree $t_{e}$ as described above. Fix a $z\in{\mathbb{R}}^{d}$, a measure $\mu\in\mathcal{P}({\mathbb{R}}^{d})$, and path measures on the ‘before’ and ‘after’ sections $\tilde{\mathbb{P}}^{t_{e},z}_{\leq}\in\mathcal{P}(\mathcal{C}_{\mathcal{T}}^{\leq}\cap\{X_{t_{e}}=z\})$ and $\tilde{\mathbb{P}}^{t_{e},z}_{\geq}\in\mathcal{P}(\mathcal{C}_{\mathcal{T}}^{\geq}\cap\{X_{t_{e}}=z\})$. Consider minimising ${\mathrm{KL}}(\cdot\|{\mathbb{Q}})$ over path measures $\mathbb{P}\in\mathcal{P}(\mathcal{C}_{\mathcal{T}})$ constrained to satisfy $\mathbb{P}_{t_{e}}=\mu$, $\mathbb{P}^{t_{e},z}_{\leq}=\tilde{\mathbb{P}}^{t_{e},z}_{\leq}$, and $\mathbb{P}^{t_{e},z}_{\geq}=\tilde{\mathbb{P}}^{t_{e},z}_{\geq}$. Then the objective ${\mathrm{KL}}(\cdot\|{\mathbb{Q}})$ attains its unique minimum at $\mathbb{P}^{*}(\cdot)=\int_{{\mathbb{R}}^{d}}\tilde{\mathbb{P}}^{t_{e},z}_{\leq}\otimes\tilde{\mathbb{P}}^{t_{e},z}_{\geq}\mu({\mathrm{d}}z)$.
@@
The proof of the claim uses Jensen’s inequality and proceeds exactly as the proof of Claim 2.11 in [60], with the appropriate notation changes. ∎
@@
As we are considering a Brownian reference process, the assumptions in question are that $\int\lVert x\rVert^{2}{\mathrm{d}}\mu_{i}(x)<\infty$, and ${\mathrm{H}}(\mu_{i})<\infty$ for each $i\in\mathcal{S}$. One can then verify (as in [23], Lemma D.2) that the criteria in Proposition B.3 hold by taking functions $A$ and $B$ to be quadratic, from which uniqueness and existence of the solution follow. From Proposition B.4 we have that the solution is in $\mathcal{R}_{\mathcal{S}}({\mathbb{Q}})$, and from Proposition B.5 we have that it is Markov.
We now need to show that if a measure $\mathbb{P}^{0}$ is Markov and in $\mathcal{R}_{\mathcal{S}}({\mathbb{Q}})$, and has the correct marginals $\mathbb{P}^{0}_{i}=\mu_{i}$ for $i\in\mathcal{S}$, then it is the TreeSB solution. Note first that as $\mathbb{P}^{0}$ is Markov and reciprocal, its restriction to each edge $e=(u,v)$ is also Markov and reciprocal along that edge. Thus, by Theorem 2.14 in [61] (noting that the required criterion holds for the Brownian reference measure) we have that $\frac{{\mathrm{d}}\mathbb{P}_{e}^{0}}{{\mathrm{d}}{\mathbb{Q}}_{e}}=f_{u}(X_{u})f_{v}(X_{v}),~~{\mathbb{Q}}_{e}\textrm{-a.e.}$ for some non-negative measurable functions $f_{u},f_{v}$. Recall that the path measures are a composition of the path measures along each edge according to the tree structure, so this means that $\frac{{\mathrm{d}}\mathbb{P}^{0}}{{\mathrm{d}}{\mathbb{Q}}}=\prod_{i\in\mathcal{V}}f_{i}(X_{i}),~~{\mathbb{Q}}\textrm{-a.e.}$ for some non-negative measurable functions $f_{i}$ (via relabelling of the functions). Note too that $\mathbb{P}^{0}$ is in $\mathcal{R}_{\mathcal{S}}({\mathbb{Q}})$, so we can also express the Radon-Nikodym derivative as $\frac{{\mathrm{d}}\mathbb{P}^{0}}{{\mathrm{d}}{\mathbb{Q}}}=h(\{X_{i}\}_{i\in\mathcal{S}})$ for some non-negative measurable function $h$. Equating the two expressions, we see that we in fact must have a decomposition only over vertices in $\mathcal{S}$, $\frac{{\mathrm{d}}\mathbb{P}^{0}}{{\mathrm{d}}{\mathbb{Q}}}=\frac{{\mathrm{d}}\Pi^{0}}{{\mathrm{d}}{\mathbb{Q}}_{\mathcal{V}}}=\prod_{i\in\mathcal{S}}f_{i}(X_{i}),~~{\mathbb{Q}}\textrm{-a.e.}$ for some non-negative measurable functions $f_{i}$ (where $\Pi^{0}$ denotes the static coupling of $\mathbb{P}^{0}$ over the vertices $\mathcal{V}$).
The remainder follows the standard argument characterising the SB solution using the decomposition according to potentials (see e.g. [65]). Consider static couplings in the constraint set $\Pi\in\Gamma(\{\mu_{i}\}_{i\in\mathcal{S}}):=\{\Pi\in\mathcal{P}(({\mathbb{R}}^{d})^{|\mathcal{V}|}:\Pi_{i}=\mu_{i}~~\forall i\in\mathcal{S}\}$ such that ${\mathrm{KL}}(\Pi|{\mathbb{Q}}_{\mathcal{V}})<\infty$. By the above, we have that ${\mathbb{E}}_{\Pi}[\log(\tfrac{{\mathrm{d}}\Pi^{0}}{{\mathrm{d}}{\mathbb{Q}}_{\mathcal{V}}})]=\sum_{i\in\mathcal{S}}\int\log f_{i}{\mathrm{d}}\mu_{i}$, which is in particular independent of the choice of $\Pi$ (for a precise statement taking care regarding the integrability of the potentials, follow the argument of Proposition 2.17 in [65]). Therefore, for any such $\Pi$ we have
@@ B.3 Properties of the tree-structured projections
We now move on to proving properties of the TreeIMF procedure. We follow the presentation of [77]. We begin by proving the properties of the reciprocal and Markovian projections defined in Definitions 3.3 and 3.4. For a full set of required assumptions, see the assumptions A.1, A.2 and A.3 in [77] Appendix C.2, which are standard in the literature and we assume to hold along each edge.
@@
[…]
@@
${\mathrm{KL}}\big($ $\mathbb{P}(\cdot|X^{e_{k}},...,X^{e_{1}},X_{r})\|\tilde{\mathbb{P}}(\cdot|X^{e_{k}},...,X^{e_{1}},X_{r})\big)$
@@
[…]
@@
We now consider the Markov projection defined in Definition 3.4. Note that the restriction of the TreeSB solution to each edge is itself an SB (because it is Markov and reciprocal), and thus can be associated with an SDE [22]. In the definition of the Markov class in Definition 3.1 we define the Markov class $\mathcal{M}_{\mathcal{T}}$ via considering SDEs with locally Lipschitz drifts. The restriction to locally Lipschitz drifts is a technical requirement for applying the entropic version of Girsanov’s theorem; this is standard in the literature and does not affect our methodology. Following [77], we now provide a result showing that the Markov projection also solves a minimisation problem.
@@
We now analyse the individual KL expressions ${\mathrm{KL}}(\Pi^{e}(\cdot|X_{u})\|{\mathbb{M}}^{e}(\cdot|X_{u}))$ along each edge $e=(u,v)$, using the proof techniques of Proposition 2 in [77].
In particular, applying the argument in the proof of Proposition 2 in [77], one sees that each conditional process $\Pi_{\cdot|0}^{e}$ is Markov and can be associated with $(X_{t}^{e})_{t\in[0,T^{e}]}$ given by
@@
Therefore, letting the restriction of the Markov process to edge $e$ (denoted above as ${\mathbb{M}}^{e}$) be associated with a process ${\mathrm{d}}Y_{t}^{e}=v_{e}(t,Y_{t}^{e}){\mathrm{d}}t+\sigma{\mathrm{d}}B_{t}^{e}$ such that ${\mathrm{KL}}(\Pi^{e}(\cdot|X_{u})\|{\mathbb{M}}^{e}(\cdot|X_{u}))<\infty$, with $v_{e}$ locally Lipschitz, then (using e.g. [53], Theorem 2.3) one obtains
@@
We also note that, as an instance of bridge matching, along each edge the process $\Pi_{t}^{e}$ and its corresponding Markovian projection ${\mathbb{M}}_{t}^{e,*}$ satisfy the same Fokker-Planck equation ([67], Theorem 2). Thus by the uniqueness of the solutions of the Fokker-Planck equations under A.1 and A.3 in [77] (see e.g. [7]), they share the same marginals ${\mathbb{M}}_{t}^{e,*}=\Pi_{t}^{e}$. ∎
@@ B.4 TreeIMF convergence
We follow the presentation of convergence in [77], but with the appropriate modifications to the proofs for the tree-based setting.
Lemma B.7 (Pythagorean property, compare to [77], Lemma 6).
@@
Using these expressions, by applying the same algebraic manipulations as the proof of Lemma 6 in [77] to each term in the summations, one obtains
@@
Proof of (44): The second part also follows similarly to [77], but instead conditioning on the values at $\mathcal{S}$. Let $\Pi^{*}=\operatorname{proj}_{\mathcal{R}_{\mathcal{S}}({\mathbb{Q}})}(\mathbb{P})=\mathbb{P}_{\mathcal{S}}{\mathbb{Q}}_{\cdot|\mathcal{S}}$. Using the change of measure formula for KL divergence and the fact that $\Pi$ and $\Pi^{*}$ have the same bridges conditional on $\mathcal{S}$,
@@
In light of the Pythagorean property in Lemma B.7, this follows from the same compactness argument as in Proposition 7 and Theorem 8 in [77]. ∎
@@ Appendix C Implementation details and extensions
Algorithm for barycentre setting
In Algorithm 2, we provide a simplified and more explicit version of Algorithm 1 for the case of computing Wasserstein barycentres.
Illustration of TreeDSBM
The TreeDSBM procedure for Wasserstein barycentre computation (that is, for a star-shaped tree) is illustrated in Figure 1. We also provide a diagram for a more general tree structure in Figure 4.
Input : Initial coupling $\Pi_{\mathcal{S}}^{0}$ over measures $\mu_{i}$ (e.g. independent),
Number of IMF iterations NN,
Entropic regularisation parameter $\sigma$.
Construct initial reciprocal process as $\Pi^{0}=\Pi_{\mathcal{S}}^{0}{\mathbb{Q}}_{\cdot|\mathcal{S}}$. That is,
  • Sample $Y_{\mathcal{S}}$ from the initial coupling $\Pi_{\mathcal{S}}^{0}$ over the marginals $\mu_{i}$;
  • Sample from the unknown central marginal as $Y_{0}\sim\mathcal{N}\big(\sum\lambda_{i}Y_{i},\sigma^{2}Id\big)$;
  • Training samples are obtained from Brownian bridges along each edge $i$ between $Y_{i}$ and $Y_{0}$.
    for $n\in\{0,\dotsc,N-1\}$ do
  •    Learn $2|\mathcal{E}|$ vector fields using bridge-matching loss (15) along each edge, using samples from current reciprocal process $\Pi^{n}$, to obtain Markovian process ${\mathbb{M}}^{n+1}$;
    Construct next reciprocal process $\Pi^{n+1}={\mathbb{M}}_{\mathcal{S}}^{n+1}{\mathbb{Q}}_{\cdot|\mathcal{S}}$ using samples from ${\mathbb{M}}_{\mathcal{S}}^{n+1}$. That is,
      Simulate ${\mathbb{M}}^{n+1}$ starting from a chosen root $r\in\mathcal{S}$ (one of the known measures $\mu_{i}$) to obtain samples $Y_{\mathcal{S}}$ from ${\mathbb{M}}_{\mathcal{S}}^{n+1}$;
    Sample from the unknown central marginal as $Y_{0}\sim\mathcal{N}\big(\sum\lambda_{i}Y_{i},\sigma^{2}Id\big)$;
    Training samples are obtained from Brownian bridges along each edge $i$ between $Y_{i}$ and $Y_{0}$.
  • end for
    Algorithm 2 TreeDSBM for Wasserstein barycentre computation